The MGR Window System HOWTO

Vincent Broman Draft 16 Nov 1994

Contents

1 This HOWTO 1 1.1 Archiving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Credit for the HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 What is the MGR window system? 2 2.1 Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 How do MGR, X11 and 8.5 compare? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Installing MGR 3 4 Running MGR 6

4.1 Applications not aware of MGR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 MGR Applications (clients) distributed with the server . . . . . . . . . . . . . . . . . . . . . . 7 4.3 MGR-aware clients distributed separately, see "SUPPORT" file . . . . . . . . . . . . . . . . . 11

5 Programming for MGR 12 6 More documentation 13 7 Credit for MGR 13

1 This HOWTO

                 Copyright Vincent Broman 1994.
           Permission granted to make and distribute
          verbatim (unaltered) copies for any purpose.

1.1 Archiving

This HOWTO is temporarily archived in ftp://bugs.nosc.mil/pub/Mgr/MGR-HOWTO, and more

permanently in ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO/MGR-HOWTO. A copy will appear in
ftp://archimedes.nosc.mil/pub/Mgr/MGR-HOWTO since bugs is disappearing in late 1994. In the same

directories may appear alternate formats like MGR-HOWTO.sgml or MGR-HOWTO.txt.

1.2 Credit for the HOWTO

While Vincent Broman first put together this HOWTO, much of the information and text was obtained from FAQs, READMEs, etc. written by Stephen Uhler, Michael Haardt, and other public-spirited net-persons. Email corrections and suggested changes to broman@nosc.mil.

Uhler was the main architect of MGR – see the Credit section below.

2 What is the MGR window system?

2.1 Function

MGR (ManaGeR) is a graphical window system. The MGR server provides a builtin window manager and windowed graphics terminal emulation on color and monochrome bitmap displays. MGR is controlled by mousing pop-up menus, by keyboard interaction, and by escape sequences written on pseudo-terminals by client software.

MGR provides each client window with: termcap-style terminal control functions, graphics primitives such as line and circle drawing; facilities for manipulating bitmaps, fonts, icons, and pop-up menus; commands to reshape and position windows; and a message passing facility enabling client programs to rendezvous and exchange messages. Client programs may ask to be informed when a change in the window system occurs, such as a reshaped window, a pushed mouse button, or a message sent from another client program. These changes are called events. MGR notifies a client program of an event by sending it an ASCII character string in a format specified by the client program. Existing applications can be integrated into the windowing environment without modification by having MGR imitate keystrokes in response to user defined menu selections or other events.

2.2 Requirements

MGR currently runs on Linux, Sun 3/4 workstations with SunOS, and Coherent. Various older versions of MGR run on the Macintosh, Atari ST MiNT, Xenix, 386-Minix, DEC 3100, and the 3b1 Unix-pc. Many small, industrial, real-time systems under OS9 or Lynx in Europe use Mgr for their user interface. The programming interface is implemented in C and in ELisp, although supporting clients written in other languages is quite easy.

Running MGR requires much less in resources than X, or even gcc. It does not have the user-base, software repertory, or high-level libraries of X or MS-Windows, say, but it is quite elegant and approachable.

It has been said that MGR is to X as Unix was to Multics.

2.3 How do MGR, X11 and 8.5 compare?

MGR consists of a server with builtin window manager and terminal emulator, and clients which run in this terminal emulator and use it to communicate with the server. No resource multiplexing is done.

X11 consists of a server and clients, which usually connect to the server using a socket. All user visible things like terminal emulators, window managers etc are done using clients. No resource multiplexing is done.

8.5, the Plan 9 window system, is a resource multiplexer, as each process running in a window can access /dev/bitblt, /dev/mouse and /dev/kbd in its own namespace. These are multiplexed to the /dev/bitblit, /dev/mouse and /dev/kbd in the namespace of 8.5. This approach allows one to run 8.5 in an 8.5 window, a very clean design. 8.5 further has an integrated window manager and terminal emulator.

3 Installing MGR

The latest source distribution can be FTPed from the directories ftp://bugs.nosc.mil/pub/Mgr/65 and

ftp://archimedes.nosc.mil/pub/Mgr/65.  One

used to be able to find older MGR sources at ftp://ftp.thp.uni-koeln.de/pub/linux/mgr, or alternatively on ftp://134.95.80.1/pub/thp/linux/mgr, but these may be gone. Even older versions of this distribution from Haardt can be found on tsx-11.mit.edu and elsewhere. Pre-Linux versions of MGR from Uhler and others can be found at ftp://bellcore.com/pub/mgr, although no one seems to maintain things there. MGR has been through a lot of versions and releases, but the current *Linux* version number is 0.65. This version number ought to arrive at 1.0 when stable 256-color VGA code for Linux appears. RCS version numbers have increased from Bellcore’s 4.3 up to our 4.12 now.

Required tools to build this distribution of MGR are m4 (GNU, or perhaps another supporting the - D option), make (GNU, or perhaps another supporting include) and *roff for the docs. Also sh, awk, and POSIX install. Binary distributions have not been assembled yet, so you need an ANSI C compiler environment, e.g. gcc.

A Linux installation requires Linux 0.99.10 or better, an HGC, EGA, VGA, or SVGA graphics card, and a mouse. Mouses supported are: serial Microsoft mouse, serial MouseSystems 3 and 5 byte mouse, serial MMSeries mouse, serial Logitech mouse, PS/2 mouse, or a bus mouse. The VGA 640x480 monochrome graphics mode is supported out of the box, as is 640x350 and 640x200. To run 800x600, or other modes that your BIOS can initialize and which do not require bank-switching, you need to run a small program (supplied as src/vgamisc/regs.exe) under DOS to read the VGA registers and write a header file which you place in the directory src/libbitblit/linux, so that it can be included by the vga.c file there. Some VGA cards can use 128k windows, and these can run higher monochrome resolutions.

The Linux-colorport code also runs in the standard 320x200x256 color VGA mode without difficulty, because no bank switching is required. Non-fast, but simple, bank-switching code has been added in version 0.65, and it works with a Tseng ET4000 card in 640x480x256 and 800x600x256 modes. The S3 code does not work in super VGA resolutions, yet. Supporting new super VGA cards requires writing one function to switch banks and making sure that the desired screen mode can be initialized from a register dump, possibly with hand-tweaking. The Linux color servers generally mangle the screen fonts, necessitating use of restorefont as in runx.

Suns with SunOS 4.1.2 and bwtwo, cgthree, or cgsix frame buffers are supported. Coherent installations should refer to the README.Coh file in the source distribution. Porting the latest-and-greatest MGR to another POSIX-like system which provides select() and pty’s and direct access to a bitmapped framebuffer ought to be straightforward, just implementing the libbitblit library based on the sunmono or

colorport code, say.

If you want to install everything, you need 5 MB disk space for binaries, fonts, manual pages etc. The sources are about 2 MB, plus object files during compilation.

Normally, /usr/mgr should be either the directory or a link to the directory where you install MGR stuff for runtime use. Typing

     chdir /usr/mgr; gunzip < whereveryouputit/mgrusr.tgz | tar xvf -

and optionally

     chdir /usr/mgr; gunzip < wherever/morefonts.tgz | tar xvf -

will unpack these. The source can be put anywhere, e.g. typing

     chdir /usr/src/local/mgr; gunzip < wherever/mgrsrc.tgz | tar xvf -

to unpack the sources from bugs.nosc.mil.

The source tree can be compiled from one top-level Makefile which invokes lower-level Makefiles, all of which "include" a "Configfile" at the top level. The Configfile is created by an interactive sh script named Configure, which runs m4 on a Configfile.m4. So you do something like this:

     chdir /usr/src/local/mgr
     sh ./Configure
     make first
     make depend
     make install
     make clean

It might be wise, before running make, to eyeball the Configfile generated by the Configure script, checking that it looks reasonable. (At least one m4 poops out (Sun /usr/bin/m4), creating a very short Configfile. If this happens, try editing a copy of Configfile.sun or Configfile.lx) Several flags in MGRFLAGS can be added/omitted to change some optional features in the server, viz:

-DWHO muck utmp file so "who" works

-DVI code for clicking the mouse in vi moving the cursor -DDEBUG

enable debugging output selectable with -d options. -DFASTMOUSE

XOR the mouse track -DBUCKEY

for hot-key server commands without mousing -DPRIORITY

for priority window scheduling instead of round-robin; the active window gets higher priority -DCUT

for cut/paste between windows and a global snarf buffer -DALIGN

forces window alignment for fast scrolling (monochr) -DKILL

kills windows upon tty i/o errors -DSHRINK

use only some of the screen ($MGRSIZE in environment) -DNOSTACK

don’t permit event stacking -DBELL

really ring the bell -DKBD

read mgr input from the sun kbd, instead of stdin. This permits redirection of console msgs to a window.

-DFRACCHAR

fractional character movement for proportional fonts -DXMENU

extended menu stuff (experimental) -DMOVIE

movie making extension which logs all operations to a file for later replay – not quite working under Linux

-DEMUMIDMSBUT Emulate a missing middle mouse button by chording Not all combinations of these options work on all systems.

The BITBLITFLAGS macro should contain -DBANKED if you’re trying out the super VGA color.

If a make complains about the lack of a default font.h or an icon server.h in the directory src/mgr, it means that you forgot to do this

     make depend

recently enough. C code for the static variables containing icons and fonts is generated by a translator from icon and font files.

Not all the clients are compiled and installed by the Makefiles. Clients found under src/clients having capitalized names or not compiled by the supplied Makefiles may have problems compiling and/or running, but they may be interesting to hack on. Most of the screen drivers found under the libbitblit directory are of mainly archeological interest. Grave robbing can be profitable.

At some point check that your /etc/termcap and/or terminfo file contain entries for MGR terminals such as found in the misc directory. If all your software checks $TERMCAP in the environment, this is not needed, as long as you run set termcap in each window.

MGR works better if run setuid root, because it wants to chown ptys and write in the utmp file. This helps the ify iconifier client work better and the event passing mechanism be more secure. On Linux, root permissions are required in order to do in/out on the screen device. Otherwise, you decide whether to trust it.

In versions around 0.62 there are troubles on the Sun with using the csh as the default shell. Programs seem to run in a different process group than the foreground process group of the window’s pty. There is no trouble with bash, sh, or rc. Ideas why?

4 Running MGR

The only file required in an MGR installation is the server itself. That would give you terminal emulator windows with shells running in them, but no nice clocks, extra fonts, fancy graphics, etc. Depending on options, a monochrome server needs about 200K of RAM plus dynamic space for windows, bitmaps, etc.

If /usr/mgr/bin is in your PATH, then just type "mgr" to start up. After enjoying the animated startup screen, press any key. When the hatched background and mouse pointer appear, hold down the left mouse button, highlight the "new window" menu item, and release the button. Then drag the mouse from corner to corner where you want a window to appear. The window will have your default shell running in it. Hold down the left mouse button over an existing window to see another menu for doing things to that window. The menu you saw that pops-up over the empty background includes the quit command. For people with a two button mouse: press both buttons together to emulate the missing middle button.

When trying to run MGR, if you get:

can’t find the screen make sure you have a /dev entry for your display device, e.g. on a Sun /dev/bwtwo0. If not, as root cd to /dev, and type "MAKEDEV bwtwo0". Otherwise, you might need the -S/dev/bwtwo0 or (on Linux) the -S640x480 command line option when starting mgr.

can’t find the mouse make sure /dev/mouse exists, usually as a symbolic link to the real device name for your mouse. If you haven’t permission to write in /dev, then something like a -m/dev/cua0 option can be given when starting mgr. Also, make sure you’ve supplied the right mouse protocol choice when you configured mgr. The mouse may speak Microsoft, even if that is not the brand name.

can’t get a pty make sure all of /dev/[tp]ty[pq]? are owned by root, mode 666, and all programs referenced with the "shell" option in your .mgrc startup file (if any) exist and are executable.

none but the default font make sure MGR is looking in the right place for its fonts. Check the Configfile in the source or see whether a -f/usr/mgr/font option to mgr fixes the problem.

completely hung (not even the mouse track moves)

login to your machine from another terminal (or rlogin) and kill the mgr process. A buckey-Q key can quit MGR if the keyboard still works.

4.1 Applications not aware of MGR

Any tty-oriented application can be run in an MGR window without further ado. Screen-oriented applications using termcap or curses can get the correct number of lines and columns by your using shape(1) to reshape the window or using set termcap(1) to obtain the correct termcap.

4.2 MGR Applications (clients) distributed with the server

bdftomgr converts some BDF fonts to MGR fonts

browse an icon browser

bury bury this window

c menu vi menus from C compiler errors

clock digital display of time of day clock2

analog display of time of day close

close this window, iconify color

set the foreground and background color for text in this window colormap

read or write in the color lookup table cursor

change appearance of the character cursor cut

cut text from this window into the cut buffer cycle

display a sequence of icons dmgr

crude ditroff previewer fade

fade a home movie script from one scene to another font

change to a new font in this window hpmgr

hp 2621 terminal emulator ico

animate an icosahedron or other polyhedron iconmail

notification of mail arrival iconmsgs

message arrival notification ify

iconify and deiconify windows loadfont

load a font from the file system maze

a maze game mclock

micky mouse clock menu

create or select a pop-up menu mgr

bellcore window system server and window manager mgrbd

boulder-dash game mgrbiff

watch mailbox for mail and notify mgrload

graph of system load average mgrlock

lock the console mgrlogin

graphical login controller mgrmag

magnify a part of the screen, optionally dump to file mgrmail

notification of mail arrival mgrmode

set or clear window modes mgrmsgs

message arrival notification mgrplot

Unix "plot" graphics filter mgrsclock

sandclock mgrshowfont

browse through mgr fonts mgrsketch

a sketching/drawing program mgrview

view mgr bitmap images mless

start up less/more in separate window, menu added for less mphoon

display the current phase of the moon mvi

start up vi in a separate window, mouse pointing oclose

(old) close a window omgrmail

(old) notification of mail arrival pbmrawtomgr

convert pbm raw bitmap to mgr bitmap format pbmstream

split out a stream of bitmaps pbmtoprt

printer output from PBM pgs

ghostscript patch and front end pilot

a bitmap browser resetwin

cleanup window state after client crashes messily rotate

rotate a bitmap 90 degrees. screendump

write graphics screen dump to a bitmap file set console

redirect console messages to this window set termcap

output an appropriate TERM and TERMCAP setting setname

name a window, for messages and iconifying shape

reshape this window square

square this window squeeze

compress mgr bitmap using run-length encoding startup

produce a skeleton startup file for current window layout texmgr

TeX dvi file previewer text2font, font2text

convert between mgr font format and text dump unsqueeze

uncompress mgr bitmap using run length encoding window print

print an image of a window zoom

an icon editor bounce, grav, grid, hilbert, mgreyes, stringart, walk

graphics demos

4.3 MGR-aware clients distributed separately, see "SUPPORT" file

calctool

on-screen calculator chess

     frontend to /usr/games/chess

gnu emacs editor with lisp/term/mgr.el mouse & menu support gnuplot

universal scientific data plotting metafont

font design and creation origami

folding editor pbmplus

portable bitmap format conversions, manipulations plplot

slick scientific data plotting ?

a groff PBM driver using Hershey fonts

5 Programming for MGR

The MGR programmers manual, the C language applications interface, is found in the doc directory in troff/nroff form. It covers general concepts, the function/macro calls controlling the server, a sample application, with an index and glossary. Porting client code used with older versions of MGR sometimes requires the substitution of

         #include <mgr/mgr.h>

for

         #include <term.h>
         #include <dump.h>

and clients using old-style B XOR, B CLEAR, et al instead of BIT XOR, BIT CLR, et al can be accommodated by writing

         #define OLDMGRBITOPS
         #include <mgr/mgr.h>

Compiling client code generally requires compiler options like the following.

         -I/usr/mgr/include   -L/usr/mgr/lib -lmgr

One can get some interactive feel for the MGR server functions by reading and experimenting with the mgr.el terminal driver for GNU Emacs which implements the MGR interface library in ELisp.

The usual method of inquiring state from the server has the potential of stumbling on a race condition if the client also expects a large volume of event notifications. The problem arises if an (asynchronous) event notification arrives when a (synchronous) inquiry response was expected. If this arises in practice (unusual) then the MGR state inquiry functions would have to be integrated with your event handling loop.

The only major drawing function missing from the MGR protocol, it seems, is an area fill for areas other than upright rectangles. There is new code for manipulating the global colormap, as well as (advisory) allocation and freeing of color indices owned by windows.

If you are thinking of hacking on the server, you can find the mouse driver in mouse.* and mouse get.*, the grotty parts of the keyboard interface in kbd.c, and the interface to the display in the src/libbitblit/* directories. The main procedure, much initialization, and the top level input loop are in mgr.c, and the interpretation of escape sequences is in put window.c.

6 More documentation

The programmer’s manual is essential for concepts.

Nearly all the clients supplied come with a man page which is installed into /usr/mgr/man/man1 or man6. Other useful man pages are bitblit.3, font.5, and bitmap.5. There is some ambiguity in the docs in distinguishing the internal bitmap format found in your frame-buffer and the external bitmap format found in files, e.g. icons.

The mgr.1 man page covers command line options, commands in the ~/.mgrc startup file, mouse and menu interaction with the server, and hot-key shortcuts available on systems with such hot-keys.

Many of the fonts in /usr/mgr/font/* are described to some extent in /usr/mgr/font/*.txt, e.g. /usr/mgr/font/FONTDIR.txt gives X-style font descriptions for the fonts obtained in .bdf format. Font names end in WxH, where W and H are the decimal width and height in pixels of each character box.

7 Credit for MGR

Stephen Uhler, with others working at Bellcore, was the original designer and implementer of MGR, so Bellcore has copyrighted much of the code and documentation for MGR under the following conditions.

    * Permission is granted to copy or use this program, EXCEPT that it
    * may not be sold for profit, the copyright notice must be reproduced
    * on copies, and credit should be given to Bellcore where it is due.

One required showing of the copyright notice is the startup title screen.

Other credits to:

• Stephen Hawley for his wonderful icons.

• Tommy Frandsen for the VGA linux library.

• Tom Heller for his Gasblit library.

• Andrew Haylett for the Mouse driver code.

• Dan McCrackin for his gasblit-¿linux patches.

• Dave Gymer, dgp@cs.nott.ac.uk, for the Startrek effect fix.

• Alex Liu for first releasing a working Linux version of MGR.

• Lars Aronsson (aronsson@lysator.liu.se) for text2font and an ISO8859-1 8-bit font.

• Harry Pulley (hcpiv@grumpy.cis.uoguelph.ca, hcpiv@snowhite.cis.uoguelph.ca) for the Coherent port.

• Vance Petree & Grant Edwards & Udo Munk for their work on Hercules.

• Udo Munk for his work on serial mouse initialization & select.

• Norman Bartek & Hal Snyder at Mark Williams Co. for their help with some bugs & with Coherent device drivers.

• Extra thanks to Zeyd Ben Halim for lots of helpful patches, especially the adaptation of selection.

• Bradley Bosch, brad@lachman.com, for lots of patches from his 3b1 port, which fix bugs and implement new and desirable features.

• Andrew Morton, applix@runxtsa.runx.oz.au, who first wrote the cut-word code.

• Kapil Paranjape, kapil@motive.math.tifr.res.in, for the EGA support.

• Michael Haardt for MOVIE support fixes, bug fixes, separation of the libbitblit code into output drivers, expansion of the libmgr, and origami folding of the code.

• Yossi Gil for many fonts.

• Carsten Emde, carsten@thlmak.pr.net.ch, for mphoon.

• Vincent Broman for middle mouse-button emulation, linting, Sun cgsix support, VGA colormap acess,

and integration of the sunport code into Haardt’s layering scheme. All bitmap fonts from any source are strictly public domain in the USA. The 575 fixed-width fonts supplied with MGR were obtained from Uhler, the X distribution, Yossi Gil, and elsewhere. The Hershey vector fonts and the code for rendering them are probably freely redistributable.